OpenStudy (anonymous):

What are the next four terms in the following geometric sequence? 5, 15, 45... A. 225; 1,125; 5,625; 28; 125 B. 135; 405; 1,215; 3,645 C. 90, 180, 360, 720 D. 90, 135, 180, 225

OpenStudy (anonymous):

geometric means you multiply on term by the same number each time to get the next one in your case that number is 3, because \(3\times 5=15\) and \(3\times 15=45\)

OpenStudy (anonymous):

so you want to answer by computing \(3\times 45\) for the next number

OpenStudy (anonymous):

@Nicin1157 you got this?